OpenStudy (anonymous):

PLEASE HELP! CALCULUS QUESTION! A company's profit from producing x tons of polyurethane is P(x)=square root of x^3-3x+34 thousand dollars for 0

OpenStudy (immanuelv):

is that from chapter 2?

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

I don't understand what the question is asking. It gave us the function for profit, and I'm assuming the derivative of the profit function is the marginal profit function. Alright, but what exactly is the answer we are looking for? I don't get it.

OpenStudy (anonymous):

well we have to find the marginal profit earned by the 5th ton and i have the answer which is 2.43 I just need the solution like how did u get 2.43

OpenStudy (anonymous):

Oh. Well I think the way you find the answer is take the derivative of the profit function (use the chain rule). Then plug in 5 as x into the derivative and you should get 2.43 if your answer is correct.

OpenStudy (anonymous):

can you show me the solution

OpenStudy (anonymous):

So this is the equation P(x)? \[\sqrt{x^3 -3x+34}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay well when I took the derivative I got this \[\frac{ 3x^2 -3 }{ 2\sqrt{x^3 -3x+34} }\]

OpenStudy (anonymous):

Now plug in 5 into x in this new equation. Problem is when I do it I got this to equal 3 not 2.43. So you sure your answer is correct?

OpenStudy (anonymous):

yeah its on my exam review professor gave the answer

OpenStudy (anonymous):

Hmm, well let me ask you this are you taking a differential calculus class or integral calculus? Is this calc 1 or calc 2 class?

OpenStudy (anonymous):

intro to calc

OpenStudy (anonymous):

Yeah well idk sorry, my calculator which can calculate derivatives also gets the same answers I do. So either I'm not understanding the question or I think ur answer sheet is wrong.

OpenStudy (anonymous):

oh okay anyways thank you for your time and do you know what is break even point?

OpenStudy (anonymous):

I think break even is when profit = 0. So take the original equation (Not the derivative) and find the x point at which the equation = 0.